# Choked precessing jets in tidal disruption events and high-energy neutrinos

### v1

@@ 3.1 Could the neutrinos from AT2019dsg, AT2019fdr and AT2019aalc be produced by choked precessing jets?
The mass of the disrupted stars in three TDEs is not well-known. Similar to Zheng et al. (2023), we assume the mass of the disrupted stars scales with the bolometric energy, i.e., $M_{\star}\approx E_{\rm bol}\approx L_{\rm acc,max}\tau$ (Metzger & Stone, 2016). Then the masses of the disrupted stars are estimated to be $0.8M_{\odot}$, $20M_{\odot}$ and $4M_{\odot}$ for AT2019dsg, AT2019fdr and AT2019aalc, respectively. Another important parameter is the mass-loss rate $\dot{M}_{\rm w}=f_{\rm w}\dot{M}$ in the disk wind, where $\dot{M}=M_{\star}/(3\tau)$ is the mass fallback rate. The characteristic fallback timescale $\tau$ for three TDE AT2019dsg, AT2019fdr, AT2019aalc are 40 days, 180 days and 140 days respectively (van Velzen et al., 2021).
@@
The expected neutrino numbers of three TDEs are shown by the color intensity in Fig.5. The dashed white lines show the parameter values corresponding to the minimum neutrino number of each TDE required by observations at $90\%$ confidence(Stein et al., 2021). The solid black lines show the critical jet luminosity limited by the break-out condition and the solid white lines show the jet luminosity limited by the power of the disk wind. Fig.5 shows that precessing choked jet model can explain the neutrinos from AT2019aalc, while it can only marginally explain the neutrinos from AT2019dsg and AT2019fdr for a fixed large misaligned angle $\theta_{\rm LS}=\pi/4$. Fig.6 illustrates the impact of the misaligned angle $\theta_{\rm LS}$ on the expected neutrino number. A weakly misaligned angle can result in a high duty cycle, consequently increase the neutrino flux. For AT2019dsg and AT2019fdr, misaligned angles less than 0.4 radians can meet the requirement of the observed neutrino flux. The most optimistic estimate of the neutrino number in the weakly misaligned jet scenario can reach $N_{\nu}=0.02$, $N_{\nu}=0.015$ and $N_{\nu}=0.12$ for AT2019dsg, AT2019fdr and AT2019aalc respectively.
@@ 3.2 Diffuse neutrino emission from Choked TDE jets
where $E_{\nu}=f_{\rm p\gamma}E_{\rm p}/8\ln(\epsilon_{\rm p,max}/\epsilon_{\rm p,min})$ is the muon neutrino energy released by a single TDE event, $E_{\rm p}$ is total CR proton energy injected by one TDE event, $\rho_{0}$ is the local event rate of TDEs, $H_{0}$ is the Hubble constant, and $f_{\rm z}$ is the correction factor for the contribution from high redshift TDEs, which is uncertain from current observations. For the TDE event rates, a latest demography study suggested that the local TDE event rate is $\rho_{0}\simeq 310{\rm Gpc^{-3}\,yr^{-1}}$ (Yao et al., 2023). A typical solar mass TDE can eject half of the mass into the disk wind at a velocity $v_{\rm w}\simeq 0.1{\rm c}$(Metzger & Stone, 2016), leading to a kinetic energy of $E_{\rm k}\approx 2.3\times 10^{51}(f_{w}/0.5)(v_{\rm w}/0.1{\rm c})^{2}$ erg in the disk wind. Noting that this energy does not violate the constraints from radio observations of some TDEs, as the ejecta is still in the free-expansion phase(Matsumoto & Piran, 2021). Radio observations show that the kinetic energy of the jets in Swift J1644+57 and AT2022cmc is about $E_{\rm k}\sim 10^{51}-10^{52}{\rm erg}$ (Cendes et al., 2023). This motivated us to assume that the kinetic energy of the choked jet is comparable to the kinetic energy of the disk wind, so the CR proton energy in one TDE event is $E_{\rm p}\sim 4.5\times 10^{50}(v_{\rm w}/0.1{\rm c})^{2}$ erg if taking $\epsilon_{\rm p}=0.2$. Assuming $f_{\rm p\gamma}\simeq 1$ and $f_{\rm z}\simeq 1$, the single flavor neutrino flux contributed by choked TDE jets is approximately $\sim 10^{-9}{\rm GeV\,cm^{-2}\,s^{-1}\,sr^{-1}}$, which is about $\sim 10\%$ of the diffuse neutrino flux at 30-300 TeV (Abbasi et al., 2022).
@@ 4 Summary and Discussions
We revisited the jet break-out condition for misaligned precessing jets in TDEs by considering that the jet could be collimated by the cocoon pressure while the jet is propagating in the quasi-isotropic disk wind. We find that the collimation effect is important for choked precessing jets. As a result, the jet head opening angle decreases as the jet propagates in the envelope, and the isotropic jet luminosity $L_{\rm j,iso}$ increases correspondingly. Nevertheless, the break-out condition does not change if we consider the beam-corrected jet luminosity, which is simply $L_{\rm j}\geq\frac{1}{2}\dot{M}f_{\rm w}c^{2}\beta_{\rm w}\pi^{2}{\rm sin}^{2}\theta_{\rm LS}$. Thus, for known values of $f_{\rm w}$, $\beta_{\rm w}$ and $\theta_{\rm LS}$, the break-out condition depends only on the jet efficiency $\eta=L_{\rm j}/\dot{M}c^{2}$. For typical parameter values of $f_{\rm w}=0.5$, $\beta_{\rm w}=0.1$ and $\theta_{\rm LS}=\pi/4$, successful jets imply $\eta>10^{-1}$. Such high-efficiency jets may not be common in TDEs, so the majority of TDE jets are choked, explaining the small fraction of TDEs that have powerful jets. At later times, the jets could possibly break out if the disk finally aligns itself with the spin axis due to the damping of the precession (Teboul & Metzger, 2023; Lu et al., 2023).
@@
We thank Wenbing Lu and Weihua Lei for valuable discussions. The work is supported by the NSFC under grants Nos. 12333006, 12121003 and 12393852.

### v3

@@ 3.1 Could the neutrinos from AT2019dsg, AT2019fdr and AT2019aalc be produced by choked precessing jets?
The mass of the disrupted stars in three TDEs is not well-known. Similar to Zheng et al. (2023), we assume the mass of the disrupted stars scales with the bolometric energy, i.e., $E_{\rm bol}\propto M_{\star}$ (Metzger & Stone, 2016). Then the masses of the disrupted stars are estimated to be $0.8M_{\odot}$, $20M_{\odot}$ and $4M_{\odot}$ for AT2019dsg, AT2019fdr and AT2019aalc, respectively. Another important parameter is the mass-loss rate $\dot{M}_{\rm w}=f_{\rm w}\dot{M}$ in the disk wind, where $\dot{M}=M_{\star}/(3\tau)$ is the mass fallback rate. The characteristic fallback timescale $\tau$ for three TDE AT2019dsg, AT2019fdr, AT2019aalc are 40 days (van Velzen et al., 2021), 180 days and 140 days calculated by Eq.(12) respectively.
@@
The expected neutrino numbers of three TDEs are shown by the color intensity in Fig.5. The dashed white lines show the parameter values corresponding to the minimum neutrino number of each TDE required by observations at $90\%$ confidence(Stein et al., 2021). The solid black lines show the critical jet efficiency limited by the break-out condition and the solid white lines show the jet efficiency limited by the power of the disk wind. The break-out criterion is normalized based on jet efficiency as $\eta\geq\frac{1}{2}f_{\rm w}\beta_{\rm w}\pi^{2}{\rm sin}^{2}\theta_{\rm LS}$. Meanwhile the critical jet efficiency $\eta_{c}=\frac{1}{2}f_{\rm w}\beta_{\rm w}^{2}=0.0025$. Fig.5 shows that precessing choked jet model can explain the neutrinos from AT2019aalc, while it can only marginally explain the neutrinos from AT2019dsg and AT2019fdr for a fixed large misaligned angle $\theta_{\rm LS}=\pi/4$. Fig.6 illustrates the impact of the misaligned angle $\theta_{\rm LS}$ on the expected neutrino number. A weakly misaligned angle can result in a high duty cycle, consequently increase the neutrino flux. For AT2019dsg and AT2019fdr, misaligned angles less than 0.4 radians can meet the requirement of the observed neutrino flux. The most optimistic estimate of the neutrino number in the weakly misaligned jet scenario can reach $N_{\nu}=0.02$, $N_{\nu}=0.015$ and $N_{\nu}=0.12$ for AT2019dsg, AT2019fdr and AT2019aalc respectively.
@@ 3.2 Diffuse neutrino emission from Choked TDE jets
where $E_{\nu}=f_{\rm p\gamma}E_{\rm p}/8\ln(\epsilon_{\rm p,max}/\epsilon_{\rm p,min})$ is the muon neutrino energy released by a single TDE event, $E_{\rm p}$ is total CR proton energy injected by one TDE event, $\rho_{0}$ is the local event rate of TDEs, $H_{0}$ is the Hubble constant, and $f_{\rm z}$ is the correction factor for the contribution from high redshift TDEs, which is uncertain from current observations. For the TDE event rates, a latest demography study suggested that the local TDE event rate is $\rho_{0}\simeq 310{\rm Gpc^{-3}\,yr^{-1}}$ (Yao et al., 2023). A typical solar mass TDE can eject half of the mass into the disk wind at a velocity $v_{\rm w}\simeq 0.1{\rm c}$(Metzger & Stone, 2016), leading to a kinetic energy of $E_{\rm k}\approx 2.3\times 10^{51}(f_{\rm w}/0.5)(v_{\rm w}/0.1{\rm c})^{2}$ erg in the disk wind. Noting that this energy does not violate the constraints from radio observations of some TDEs, as the ejecta is still in the free-expansion phase(Matsumoto & Piran, 2021b). Radio observations show that the kinetic energy of the jets in Swift J1644+57 and AT2022cmc is about $E_{\rm k}\sim 10^{51}-10^{52}{\rm erg}$ (Cendes et al., 2023). This motivated us to assume that the kinetic energy of the choked jet is comparable to the kinetic energy of the disk wind, so the CR proton energy in one TDE event is $E_{\rm p}\sim 4.5\times 10^{50}(v_{\rm w}/0.1{\rm c})^{2}$ erg if taking $\epsilon_{\rm p}=0.2$. Assuming $f_{\rm p\gamma}\simeq 1$ and $f_{\rm z}\simeq 1$, the single flavor neutrino flux contributed by choked TDE jets is approximately $\sim 10^{-9}{\rm GeV\,cm^{-2}\,s^{-1}\,sr^{-1}}$, which is about $\sim 10\%$ of the diffuse neutrino flux at 30-300 TeV (Abbasi et al., 2022).
@@ 4 Summary and Discussions
We revisited the jet break-out condition for misaligned precessing jets in TDEs by considering that the jet could be collimated by the cocoon pressure while the jet is propagating in the quasi-isotropic disk wind. We find that the collimation effect is important for choked precessing jets. As a result, the jet head opening angle decreases as the jet propagates in the envelope, and the isotropic jet luminosity $L_{\rm j,iso}$ increases correspondingly. Nevertheless, the break-out condition does not change if we consider the beam-corrected jet luminosity, which is simply $L_{\rm j}\geq\frac{1}{2}\dot{M}f_{\rm w}c^{2}\beta_{\rm w}\pi^{2}{\rm sin}^{2}\theta_{\rm LS}$. Thus, for known values of $f_{\rm w}$, $\beta_{\rm w}$ and $\theta_{\rm LS}$, the break-out condition depends only on the jet efficiency $\eta=L_{\rm j}/\dot{M}c^{2}$. For typical parameter values of $f_{\rm w}=0.5$, $\beta_{\rm w}=0.1$ and $\theta_{\rm LS}=\pi/4$, successful jets imply $\eta>10^{-1}$.
An important result of Lu et al. (2023) is that a successful jet breakout requires a double alignment condition, which naturally explains why the event rate of jetted TDEs is much smaller than optical/X-ray TDEs. They find that the fraction of jetted TDEs that meet the double alignment condition is $\sim 10^{-3}({\theta_{j}}/0.1)^{4}$ assuming a half jet opening angle of $\theta_{j}=0.1$. The $\theta_{j}=0.1$ case of a conical jet in Lu et al. (2023) corresponds to $\theta_{0}\sim 0.33$ in our case, which has a half jet opening angle $0.1$ after collimation.
@@
If the jet luminosity is limited by the wind power, a majority of TDE jets may have lower efficiency than observed jetted TDEs. We assume the critical value $\eta_{c}$ as a typical choked jet efficiency, then the dimensionless black hole spin parameter is $a\sim 0.2-0.3$ by utilizing $\eta=4\times 10^{-3}\Phi_{\rm B}^{2}\Omega_{\rm H}^{2}(1+1.38\Omega_{\rm H}^{2}-9.2\Omega_{H}^{4})$, where $\Omega_{\rm H}=a/(1+(1-a^{2})^{1/2})$ is the dimensionless angular frequency of the horizon and $\Phi_{\rm B}\simeq-20.2a^{3}-14.9a^{2}+34a+52.6$ is dimensionless magnetic flux threading the black hole’s horizon(Teboul & Metzger, 2023; Lu et al., 2023).
@@
We thank Wenbin Lu and Weihua Lei for valuable discussions. The work is supported by the NSFC under grants Nos. 12333006, 12121003 and 12393852.